OpenStudy (anonymous):

solve this system of equation using subtraction 8x - y = 4 5x + 3y = 17

OpenStudy (ganpat):

you can do this.. right ?? refer to previous problem..

OpenStudy (anonymous):

tbh no

OpenStudy (ganpat):

ohkey understand this... Multiply equation 1 by -3... (why -3 ?? because we can eliminate one variable..) so -24x + 3y = -12... eqtn 1 5x + 3y = 17 ..... equation 2.. Now subtracting both of them.. -29x = -29.. so x = 1... substitute the value in any of the equation above.. 5(1) + 3y = 17 3y = 12 y = 4.. Now you understand, and u should know... right ??

OpenStudy (anonymous):

according to the problem x=1 y=4 is wrong

OpenStudy (goformit100):

YES

OpenStudy (rishabh.mission):

x=1 and y=4 is true.

OpenStudy (ganpat):

@goformit100 : right answer ??